XYZ is a triangle with side lengths in the ratio 2 : 3 : 4. The perimeter of XYZ is 27 yards. Find the length of the longest side.
Answer by mananth(16949)

ratios 2:3:4
multiply by common factor x
2x , 3x, 4x are the lenngths
2x+3x+4x = 27
9x=27
/9
x=3
longest side = 4x ==> 4*3 =12 yards